# On the effectiveness of discrete representations in sparse mixture of experts

### v1

@@ Appendix B Proof for Results in Section 4
According to (Chi et al., 2022), projecting information from a hidden representation space $\mathcal{R}^{d}$ to the expert dimension $N$ leads to representation collapse. Now, consider three experts $x,y,z$ whose embeddings $e_{x},e_{y},e_{z}$ collapse. Without loss of generality, assume that $e_{y}$ lies between $e_{x}$ and $e_{z}$ in the embedding space. Then, we have:
$\text{dist}(y,u_{y})\leq\min(\text{dist}(x,e_{x}),\text{dist}(y,e_{y}),\text{dist}(z,e_{z}))\leq\text{dist}(e_{x},e_{z}).$
Let $t_{e}$ denote the step at which the embeddings $e_{x}$ and $e_{z}$ converge, and $t_{m}$ denote the step at which the Multi-Head Attention (MHA) module converges. From step $t_{e}$, it follows that:
$\lim_{t_{e}\to t_{m}}\text{dist}(y,u_{y})=\lim_{t_{e}\to t_{m}}\text{dist}(e_{x},e_{z})=0.$
@@
B.2 Proof of Proposition 4.2
We use contradiction to prove the proposition. Assume that, at training step $t$, there exists a set of pairs $(C_{i},E_{j})$ such that $i\neq j$. Let $x_{1},x_{2},\ldots,x_{k}$ represent a sequence of inputs sampled from $K$ clusters. From step $t_{0}$ to step $t_{k-1}$, each pair $(x_{j},E_{j})$, where $j\in[1,k]$, is updated using the following gradient descent equation:
$W^{l+1}_{E_{j}}=W^{l}_{E_{j}}-\eta\mathcal{J}(x_{j}),$
where $W^{l}_{E_{j}}$ is the weight of expert $E_{j}$ at iteration $l$, $\mathcal{J}(x_{j})$ is the Jacobian matrix with respect to input $x_{j}$, and $\eta$ is the learning rate.
Let $\mathcal{L}$ denote the loss function during the training process described by Equation 6. After $t_{k}$ training steps, the following condition holds:
$E_{j}(x_{j})=\min_{c\in[1,k]}E_{j}(x_{c}).$
Under the assumption of contradiction, there exists a set of pairs
$\sum_{i,j=1;i\neq j}^{K}(C_{i},E_{j})$
where the loss function $\mathcal{L}$ is minimized. However, by definition of the loss minimization process, the inequality
$\sum_{i=1}^{K}(C_{i},E_{i})\leq\sum_{i,j=1;i\neq j}^{K}(C_{i},E_{j})$
@@
Appendix C Representation Collapse Analysis
@@
To illustrate Theorem  4.1, we perform a language model task as described in Section  D.2, examining the movement of Expert Input Representation in Figure  4 and Expert Embedding (router) in Figure  5. We analyze the dynamics of the expert input representations by tracking their changes across training iterations. The results indicate that the inputs to the experts become increasingly divergent over time. This divergence suggests that the model learns to represent the data in a more specialized and diverse manner, allowing each expert to focus on distinct features or patterns within the data. Similarly, we track the changes in expert embeddings (router) throughout the training process. However, the trend is the opposite: the expert embeddings appear to converge quickly, stabilizing around 10,000 iterations. The findings align with our assumption stated in Theorem  4.1, indicating that Expert Embedding converges more quickly than Expert Input Representation. These results provide further evidence supporting the Theorem  4.1.
Figure 4: Training SMoE Expert Input Representations across Training Iterations.
Figure 5: Training SMoE Router (Expert embedding) across Training Iterations.
(a) Vector Quantization method.
(b) Number of codebook.
(c) Impact of $\alpha$ for VQMoE.
Appendix D Experiments implementation details
@@ D.2 Pre-training Experiments
Dataset Input length Batch size Optimizer Lr # Training Step
Enwik8 512 48 Adam 3.5e-4 100k
Text 512 48 Adam 3.5e-4 100k
WikiText-103 512 22 Adam 3.5e-4 100k
One Billion Word 512 11 Adam 3.5e-4 100k
Table 4: Hyperparameter settings for pre-training experiments on Enwik8, Text8. , WikiText-103. , and One Billion Word.
D.3 Fine-tuning Experiments

### v2

@@ A.1.1 Jacobian Matrix of VQMoE
According to (Chi et al., 2022), projecting information from a hidden representation space $\mathbb{R}^{d}$ to the expert dimension $N$ leads to representation collapse. Now, consider three output of Multi-Head Attention (MHA) (MHA) layer: $x_{1},x_{2},x_{3}$ $\in\mathbb{R}^{d}$, belong to experts whose embeddings $e_{1},e_{2},e_{3}$ collapse. Without loss of generality, assume that $e_{2}$ lies between $e_{1}$ and $e_{3}$ in the embedding space. Then, we have:
$\text{dist}(x_{2},u_{2})$ $\leq\min(\text{dist}(x_{1},e_{1}),\text{dist}(x_{2},e_{2}),\text{dist}(x_{3},e_{3}))$ (13)
$\leq\text{dist}(e_{1},e_{3}).$
Let $t_{e}$ denote the step at which the embeddings $e_{1}$ and $e_{3}$ converge, and $t_{m}$ denote the step at which the Multi-Head Attention (MHA) module converges. From step $t_{e}$, it follows that:
$\lim_{t_{e}\to t_{m}}\text{dist}(x_{2},u_{2})=\lim_{t_{e}\to t_{m}}\text{dist}(e_{1},e_{3})=0.$
@@
A.1.3 Proof of Proposition 4.4
We use contradiction to prove the proposition. Assume that, at training step $t$, there exists a set of pairs $(C_{i},E_{j})$ such that $i\neq j$. Let $x_{1},x_{2},\ldots,x_{N}$ represent a sequence of inputs sampled from $N$ clusters. From step $t_{0}$ to step $t_{m-1}$, each pair $(x_{j},E_{j})$, where $j\in[1,N]$, is updated using the following gradient descent equation:
$W^{t_{l+1}}_{E_{j}}=W^{t_{l}}_{E_{j}}-\eta\mathcal{J}(x_{j}),$
where $W^{t_{l}}_{E_{j}}$ is the weight of expert $E_{j}$ at iteration $t_{l}$, $\mathcal{J}(x_{j})$ is the Jacobian matrix with respect to input $x_{j}$, and $\eta$ is the learning rate, and $0\leq l<m-1$.
Let $\mathcal{L}$ denote the loss function during the training process described by Equation 8. After $t_{m-1}$ training steps, the following condition holds:
$\mathcal{L}(E_{j}(x_{j}))=\min_{c\in[1,N]}\mathcal{L}(E_{c}(x_{j})).$
Under the assumption of contradiction, there exists a set of pairs, where $x_{j}$ is assigned to an expert $E_{i}$: $(x_{j},E_{i})$ ; $i,j\in[1,N]$ and $i\neq j$; where the loss function $\mathcal{L}$ is minimized. It means:
$\mathcal{L}(E_{i}(x_{j}))\leq\mathcal{L}(E_{j}(x_{j}))$
However, by definition of the loss minimization process, the inequality
$\mathcal{L}(E_{j}(x_{j}))\leq\mathcal{L}(E_{i}(x_{j}))$
@@
A.2 Additional Experiment Results
Q6: Can VQMoE learn Discrete Representation Only from scratch? A6: Yes for small scale, but no for large scale.
The answer is yes for small models. However, training a discrete representation-only approach is feasible primarily for small-scale models with a moderately sized dataset. The results of the Transformer-XL model in Table 4 on the Enwik8 dataset support this observation. As the model scales up, relying solely on discrete representation reaches its limitations, leading to performance below the SMoE baselines.
Scale TopK # Experts SMoE VQMoE (Discrete Only)
Base 20M-50K Steps 1 16 1.28 1.25
2 16 1.26 -
4 16 1.26 -
8 16 1.27 -
16 16 1.27 -
Base 20M-100K Steps 1 16 1.22 1.18
2 16 1.20 -
4 16 1.21 -
8 16 1.21 -
16 16 1.21 -
Large (210M) 1 64 1.12 1.14
2 64 1.09 -
4 64 1.09 -
8 64 1.09 -
16 64 1.10 -
32 64 1.10 -
64 64 1.12 -
Table 4: Performance comparison of SMoE and VQMoE (Discrete Only) on the Enwik8 (BPC) dataset.
Q7: Can VQMoE outperform the clustering-based approach such as KMean? A7: Yes.
We explored a clustering-based approach -MoCLE(Gou et al., 2024), but found it unsuitable for our method. Unlike MoCLE, Vector Quantization allows the model greater flexibility in learning cluster representations during training, making it more competitive in practical applications. The training results using the Transformer-XL model on the Enwik8 dataset are presented in Table 5.
Scale TopK # Experts SMoE MoCLE VQMoE
Base 20M-50K Steps 1 16 1.28 1.29 1.25
2 16 1.26 1.28 -
4 16 1.26 1.28 -
8 16 1.27 1.28 -
16 16 1.27 1.28 -
Table 5: Performance comparison of VQMoE and MoCLE (Clustering approach) on the Enwik8 (BPC) dataset.
Q8: Can VQMoE contribute to AI real-world applications? A8: Yes.
We found that VQMoE can directly benefit real-world AI applications, such as image segmentation, demonstrating its strong generalization capabilities. Specifically, our method outperforms both the baseline and dense models in terms of Mean Accuracy and mIoU metrics on the ADE20K dataset (Zhou et al., 2018) using the Segmenter model(Strudel et al., 2021). Detailed results are provided in Table 6.
Model ViT SoftMoe SMoE StableMoE XMoE VQMoE Metrics
Segmenter 20.8 19.0 23.1 22.4 22.3 23.4 Mean accuracy
15.0 14.0 15.5 16.0 15.7 16.6 mIoU
Table 6: Comparison of VQMoE versus the baselines on the ADE20K dataset.
Q9: Does VQMoE consistently outperform the baselines across multiple training runs? A9: Yes.
Due to resource constraints, it is challenging to train all models across all datasets multiple times and to perform formal statistical significance testing. To illustrate the variance across multiple training runs, we train VQMoE and baseline models on the Text8 dataset three times. The average Bits-Per-Character (BPC) and standard deviation for each model are reported in Table 7. The results indicate that VQMoE achieves the best average performance, while also exhibiting a lower standard deviation compared to other models, suggesting greater training stability. The consistency observed across repeated runs supports the reliability of the results reported in Table 7.
Table 7: Average BPC and standard deviation across three training runs on Text8. Lower is better; best results are in bold.
Model Algorithm Dataset Avg. BPC
[…]
SMoE 1.293±0.0071.293$\pm$ 0.0071.293 ± 0.007
XMoE 1.282±0.0031.282$\pm$ 0.0031.282 ± 0.003
StableMoE 1.285±0.0051.285$\pm$ 0.0051.285 ± 0.005
Q10: Is VQMoE able to consistently surpass SMoE models in large-scale evaluation scenarios? A10: Yes.
[…]
Table 8: Zero-shot performance comparison between OLMoE and VQMoE on MTEB. The best score per dataset is highlighted in bold. Improvement (Imp.) is calculated as (GAP / OLMoE) * 100
[…]
Classification Emotion 7B 64 8 49.9 52.5 5.2
Toxic 7B 64 8 65.2 67.2 3.1
Tweet 7B 64 8 58.0 59.8 3.1
Clustering Medrxiv 7B 64 8 23.9 25.8 7.5
20Groups 7B 64 8 25.7 28.4 10.5
Pair Classification SemEval 7B 64 8 46.7 49.5 6.0
URLCorpus 7B 64 8 77.4 79.4 2.6
Reranking Ask 7B 64 8 51.9 53.3 2.7
SciDocs 7B 64 8 69.6 72.3 3.7
StackOver 7B 64 8 32.5 33.9 4.3
STS Biosses 7B 64 8 61.8 68.7 11.2
SickR 7B 64 8 65.7 66.5 1.4
STS12 7B 64 8 53.8 56.0 4.1
STS13 7B 64 8 66.5 74.0 11.3
STS14 7B 64 8 56.8 59.5 4.6
STS15 7B 64 8 69.3 71.5 3.2
STS16 7B 64 8 70.1 70.5 0.6
Summarization Medrxiv 7B 64 8 28.9 29.8 3.1
Average 54.1 56.6 4.6
Interestingly, we find that VQMoE consistently outperforms OLMoE across all tasks and datasets, despite not undergoing additional training or fine-tuning. On average, across six tasks, VQMoE shows a relative improvement of 4.6%. The most significant gains appear in the Classification and Clustering tasks. These findings support our hypothesis that VQMoE enhances pre-trained models by learning more effective routing policies. Furthermore, by mitigating representation collapse through the use of discrete representations, VQMoE improves the model’s overall representational capacity.
A.3 Representation Collapse Analysis
@@
To illustrate Theorem  4.3, we perform a language model task as described in Section  A.4.2, examining the movement of Expert Input Representation in Figure  5(a) and Expert Embedding (router) in Figure  5(b). We analyze the dynamics of the expert input representations by tracking their changes across training iterations. The results indicate that the inputs to the experts become increasingly divergent over time. This divergence suggests that the model learns to represent the data in a more specialized and diverse manner, allowing each expert to focus on distinct features or patterns within the data. Similarly, we track the changes in expert embeddings (router) throughout the training process. However, the trend is the opposite: the expert embeddings appear to converge quickly, stabilizing around 10,000 iterations. The findings align with our assumption stated in Theorem  4.3, indicating that Expert Embedding converges more quickly than Expert Input Representation. These results provide further evidence supporting the Theorem  4.3.
A.4 Experiments implementation details
@@ A.4.2 Pre-training Experiments
Dataset Input length Batch size Optimizer Lr # Training Step # Experts TopK
Enwik8 512 48 Adam 3.5e-4 100k 16 2
Text 512 48 Adam 3.5e-4 100k 16 2
WikiText-103 512 22 Adam 3.5e-4 100k 16 2
One Billion Word 512 11 Adam 3.5e-4 100k 16 2
Table 9: Hyperparameter settings for pre-training experiments on Enwik8, Text8 , WikiText-103 , and One Billion Word.
@@
A.4.3 Fine-tuning Experiments
(a) Training Input Token Representations.
(b) Training Router Representation (Expert embedding).
Figure 5: Comparison of Token Representation and Expert Representation across Training Iteration.
(a) Vector Quantization method.
(b) Number of codebook.
(c) Impact of $\alpha$ for VQMoE.